# Durham, UK Image Inpainting

### v1

@@ 6 Conclusion
In this paper, we introduce MxT, a hybrid model for iamge inpainting designed to reconstruct high-quality images with fine-grained details and spatial coherence. The propsoed Hybrid Module effectively combines transfromer and Mamba, leveraging the capacity of mamba for capturing pixel-wise long-range interaction along with the spatial perception provided by transfromer. This integration enable MxT to maintain linear computational complexity, which is particularly advantageous for handling high-resolution image. We validate MxT on the widely-used CelebA-HQ and Places2 datasets, where it demonstrated superior or comparable performance to existing state-of-the-art methods.

### v2

@@ 6 Conclusion
In this paper, we introduce M×$\times$×T, a hybrid model for image inpainting designed to reconstruct high-quality images with fine-grained details and spatial coherence. The proposed Hybrid Module effectively combines transformer and Mamba, leveraging the capacity of mamba for capturing pixel-wise long-range interaction along with the spatial perception provided by the transformer. This integration enables M×$\times$×T to maintain linear computational complexity, which is particularly advantageous for handling high-resolution images. We validate M×$\times$×T on the widely-used CelebA-HQ and Places2 datasets, where it demonstrated superior or comparable performance to existing state-of-the-art methods.